Trainingnumber of hours spent lifting weights a week Weight 96+ 11.9(Training) The data point is an individual who trains 5 hours a week and can bench 150 pounds. 6 Your answer is partially correct. Try again. (a) Find the predicted value for the data point and compute the residual. Enter the exact answers. lbs udy Predicted value- 155.5 5.5 lbs Residual= SHOW HINT LINK TO TEXT
Chapter 2, Section 6, Exercise 202 Two variables are defined, a regression equation is given, and one data point is given. Weight Training maximum weight capable of bench pressing (pounds) number of hours spent lifting weights a week Weight 96+ 11.(Training) The data point is an individual who trains 5 hours a week and can bench 150 pounds. Your answer is partially correct. Try again. (a) Find the predicted value for the data point and compute the residual. Enter the exact answers. Predicted value- 155.5 lbs
Answer #1

Solution :-

Given data :

Regression equation is

Weight = 96 + 11.9 (training)

Here, Weight = maximum weight capable of bench pressing (pounds), and Training = number of hours spent lifting weights a week.

The data point is an individual who trains 5 hours a week and can bench 150 pounds.i.e, X = 5 and y = 150.

( a ) :-

Here we need to calculate the  predicted value for the data point.

Here data point is X = 5

Then, Weight = 96 + 11.9 (training)

   Weight = 96 + 11.9 ( 5 )

   Weight = 96 + 59.5

   Weight = 155.5

the  predicted value for the data point is hat{y} = 155.5

RightarrowHere we need to calculate the residual.

Here we know the equation i.e,

e=y-hat{y}

e 150-155.5

e=-5.5

The residual is e = -5.5